OpenStudy (anonymous):

7y + 4x = 3; (-4, -7) write the equation in slope intercept form of the line that is perpendicular to the graph of each equation and passes through the given point

OpenStudy (texaschic101):

we first need to find the slope of the equation. Do you know how to do that ?

OpenStudy (anonymous):

nope

OpenStudy (texaschic101):

in y = mx + b form, the m position is the slope. 7y + 4x = 3 (put it in y = mx + b form) 7y = -4x + 3 y = -4/7x + 3/7 the slope is -4/7 you understand so far ?

OpenStudy (anonymous):

yeah sorta..

OpenStudy (texaschic101):

now we are looking for a perpendicular line, so we need the negative reciprocal of that slope. All that means is we " flip " the slope and change the sign. So the negative reciprocal of -4/7 = 7/4 (you see how I flipped the slope and changed the sign) So we will be using slope(m) = 7/4 and (-4,-7) x = -4 and y = -7 y = mx + b now sub -7 = 7/4(-4) + b -7 = -28/4 + b -7 = - 7 + b -7 + 7 = b 0 = b so your equation is : y = 7/4x + 0

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

omg i think so

OpenStudy (texaschic101):

It will take a little time to learn this, but once you get it down....and you will...it will be easy :)

OpenStudy (anonymous):

you were such a big help. i actually get it. I'm solving one right now

OpenStudy (texaschic101):

good to hear. Keep practicing and you will learn all this in no time :)

OpenStudy (anonymous):

thank you :)